Two trains leave the station at the same time, one heading east and the other west. The eastbound train travels 18 miles per hour faster than the westbound train. If the two trains are 760 miles apart after 4 hours, what is the rate of the eastbound train?
Answer by richwmiller(17219)

4x+4(x+18)=760
8x=742
x=742/8
x=86
18+86=104 mph eastbound